Solve for all real roots
(x-4)^2=36

Found 2 solutions by jim_thompson5910, MathPro:
Answer by jim_thompson5910(13794)


Start with the given equation.


Take the square root of both sides.


or Break up the "plus/minus" to form two equations.


or Take the square root of to get .


or Add to both sides.


or Combine like terms.


--------------------------------------


Answer:


So the solutions are or .

Answer by MathPro(14)

Good evening,
Your problem is: Solve for all real roots (x-4)^2=36
The first term is squared so let’s try to get rid of that by taking the square root of both sides of the equation:
(x-4)^[2(1/2)] = 36^1/2
x - 4 = 6
Let’s solve for x by adding 4 to both sides:
x – 4 + 4 = 6 + 4
x =10
Check the answer by substituting for x in the original equation:
(10-4)^2 = 36
36 = 36
The left-hand-side is equal to the right-hand-side, so the answer is correct.
Good Luck!
Selma from Rent-A-Whiz, an online tutoring and homework help website. Just visit www.rentawhiz.com to learn more. Currently, we are offering to solve one math problem for free. Just visit our website, select the homework tab, then click on the math problem submission link and quote coupon code 003 when you submit your problem.